# Position Paper Are Normalizing Flows the Key to Unlocking the Exponential Mechanism? A Path through the Accuracy-Privacy Ceiling Constraining Differentially Private ML

### v3

@@
Abstract
The state of the art and de facto standard for differentially private machine learning (ML) is differentially private stochastic gradient descent (DPSGD). Yet, the method is inherently wasteful. By adding noise to every gradient, it diminishes the overall privacy with every gradient step. Despite 15 years of fruitful research advancing the composition theorems, sub-sampling methods, and implementation techniques, adequate accuracy and privacy is often unattainable with current private ML methods. Meanwhile, the Exponential Mechanism (ExpM), designed for private optimization, has been historically sidelined from privately training modern ML algorithms primarily because ExpM requires sampling from a historically intractable density. Despite the recent discovery of Normalizing Flow models (NFs), expressive deep networks for approximating intractable distributions, ExpM remains in the background. Our position is that leveraging NFs to circumvent historic obstructions of ExpM is a potentially transformational solution for differentially private ML worth attention. We introduce a new training method, ExpM+NF, as a potential alternative to DPSGD, and we provide experiment with logistic regression and a modern deep learning model to test whether training via ExpM+NF is viable with “good” privacy parameters. Under the assumption that the NF output distribution is the ExpM distribution, we are able to achieve $\varepsilon$ a low as $1\mathrm{e}{-3}$—three orders of magnitude stronger privacy with similar accuracy. This work outlines a new avenue for advancing differentially private ML, namely discovering NF approximation guarantees. Code to be provided after review.
differential privacy, machine learning, exponential mechanism, normalizing flows, optimization
@@ Abstract
Figure 1: Figure depicts private data enclave on which outside entities would like to train a machine learning model. New technique, ExpM+NF, depicted at right, trains an auxiliary Normalizing Flow (NF) model to sample near optimal parameters from the otherwise intractable Exponential Mechanism (ExpM) density. Privacy parameters $(\varepsilon>0,\delta=0)$ are pre-specified by the private data center.
1 Introduction
@@
Differential privacy (DP) is the widely accepted standard for privacy preserving analytics and its adoption has extended into a variety of AI/ML applications (Xiong et al., 2014). Differentially private ML is currently achieved almost exclusively by a single, preeminent technique—Differentially Private Stochastic Gradient Descent (DPSGD) (Abadi et al., 2016; Ponomareva et al., 2023). DPSGD achieves a privacy guarantee by clipping and adding Gaussian noise to each gradient during SGD training, and upon the conclusion of training the many privacy guarantees (one per step) are collected into a single, overarching guarantee. Aiding DPSGD’s widespread adoption is 15 years of work on composition theorems (Dwork, 2007, 2010; Kairouz et al., 2017; Dong et al., 2021) which include methods that tighten the privacy bounds for mechanisms that utilize subsampling (e.g., ML training with batches) (Abadi et al., 2016; Mironov, 2017) and clever techniques that compose privacy curves using the privacy random variable (Gopi et al., 2021).
@@
Despite these years of effort seeking to improve the privacy bound for ML models trained with DPSGD, one can see in the conventional scenario of training ML on a single private dataset that DPSGD is wasteful by design. Privacy is expended so that all gradients may be publicized, yet all that is required is the trained model. This leads to the overall privacy guarantee worsening with each step during training, of which hundreds to tens of thousands are usually needed. Once the privacy budget is expended, training must cease, perhaps prematurely. This has led to an unusable accuracy-privacy tradeoff in practice (Kurakin, 2022; Suriyakumar et al., 2021). Secondly, by expending privacy at each gradient step, DPSGD can only provide approximate DP, i.e., $\delta>0$ (see Definition 2.1), a consequence of relying on composition theorems and, independently, of using Gaussian noise.
In this paper, we advocate for a fundamentally different approach for achieving differentially private ML leveraging the Exponential Mechanism (ExpM) (McSherry & Talwar, 2007), a well-established DP technique for optimizing a utility (e.g., an inverse loss) function. The goal is to sample from the ExpM distribution, a probability distribution over the model’s parameter space (the domain of the utility/loss function) so that near-optimal (resp., far-from-optimal) parameters have high (resp., low) likelihood. By design, sampling from this distribution produces, with high likelihood, near-optimal parameters that are furnished with a strong privacy guarantee—pre-specified privacy $\varepsilon>0$ and $\delta=0$.
While ExpM is well known, two obstructions prevented its prior use in ML applications. First, an a priori bound on sensitivity must be proven for the given utility (equiv., loss) function, which is difficult in general, e.g., see (Minami et al., 2016). Secondly, sampling from the requisite distribution is intractable except for finite or otherwise very simple sample domains, where ExpM reduces to sampling from a multinomial. Examples of using ExpM for differentially private ML in the finite setting exist, e.g.: (Friedman & Schuster, 2010) (private decision tree with only categorical features), (Kapralov & Talwar, 2013) (private PCA), and (Vinterbo, 2012) (private feature selection with only categorical variables). Notwithstanding these examples, ExpM is historically intractable in most ML applications as the optimization is of a loss function over continuous parameter space (Kasiviswanathan et al., 2011; Ji et al., 2014; Minami et al., 2016; Near & Abuah, 2022).
For the first barrier, we constrain our training to $\ell^{2}$ (equiv. MSE) loss, admitting a sensitivity bound for supervised learning tasks with bounded targets (accommodating classification tasks via binary/one-hot encoding)—Theorem 3.1. While we do not consider other loss functions, additional theorems for other loss functions could be developed and used in a similar manner, potentially with auxiliary privacy benefits.
To address the second obstruction, we leverage an auxiliary Normalizing Flow (NF) model—a deep learning model—that permits approximate sampling from the otherwise intractable ExpM density (Papamakarios et al., 2021). We note that other approximate sampling methods that leverage ML could prove just as promising. The combined technique, ExpM+NF (see Figure 1), promises a strong user-specified privacy guarantee ($\varepsilon>0,\delta=0$) with near optimal parameters. Further, by design ExpM+NF eliminates the iterative privacy loss observed in DPSGD.
Whether or not (and specifically how) one can train a desired ML model using ExpM+NF with ideally strong privacy parameter $\varepsilon$ is a research question. To justify this new approach, we implement experiments using ExpM+NF to train logistic regression (LR) and GRU-D, a deep learning model, on MIMIC-III data (Johnson et al., 2016). This exhibits that ExpM+NF can achieve the desired model performance. Assuming that the NF output is the ExpM density, our results show that ExpM+NF achieves similar accuracy with orders of magnitude better privacy (lower $\varepsilon$ with $\delta=0$) than DPSGD. If the NF furnishes an accurate approximation of the ExpM distribution, then this method substantially improves the SOTA in differentially private ML.
We do not maintain that the privacy implications of the NF approximation can be ignored. Rather, our position is that identifying strong bounds on NF approximation accuracy and their associated privacy guarantees now promises a route through the accuracy–privacy ceiling that currently constrains differentially private ML. Such an advancement would prove transformational for differentially private ML; the vast increase in the privacy budget may be used to address other hardships, such as affording hundreds, possibly thousands of hyperparmeter tuning runs (thereby increasing accuracy) while furnishing even better privacy.
2 Background
@@
2.1 Differential Privacy
@@
A mechanism $f:\Theta\times D\to\mathbb{R}$ satisfies $(\varepsilon,\delta)$-Differential Privacy or is $(\varepsilon,\delta)$-DP if and only if
$P(f(\theta,X)\in A)\leq P(f(\theta,X^{\prime})\in A)e^{\varepsilon}+\delta$
for all neighboring $X,X^{\prime}\in D,\theta\in\Theta,\varepsilon\geq 0$ and for all measurable sets $A$. When $\delta=0$, we say $f$ is $\varepsilon$-DP.
The $\varepsilon$ term is an upper bound on the privacy loss. Approximate DP, or $(\varepsilon,\delta)$-DP, provides a relaxed definition—$\delta$ is the probability that the mechanism $f$ is not guaranteed to be $\varepsilon$-DP. This “probability of failure” interpretation has let many to suggest that $\delta$ should be no larger that $1/|X|$ (Dwork & Roth, 2014).
To achieve differential privacy in practice, randomness is leveraged. Determining how much randomness to inject requires understanding how sensitive the output of the mechanism is across all neighboring datasets.
@@
The sensitivity of a mechanism $f:\Theta\times D\to\mathbb{R}$ is $\sup|f(\theta,X)-f(\theta,X^{\prime})|$ where the supremum is over neighboring $X,X^{\prime}\in D$, and all $\theta\in\Theta$. We use $s$ to denote an upper bound on sensitivity.
Definition 2.3 (Exponential Mechanism).
For data $X\in D$ and a utility function $u:\Theta\times D\to\mathbb{R}$ with sensitivity $s$, the Exponential Mechanism (ExpM) returns $\theta\in\Theta$ with likelihood $p_{ExpM}(\theta;X)\propto\exp(\varepsilon u(\theta,X)/(2s))$.
The following theorem shows ExpM satisfies $\varepsilon$-DP.
Theorem 2.4 ((McSherry & Talwar, 2007; Dwork & Roth, 2014)).
@@
The original treatment of ExpM (McSherry & Talwar, 2007) furnishes utility theorems—rigorous guarantees that bound how far ExpM samples can be from the optima.
2.2 Normalizing Flows (NFs)
We reference readers to the survey of NF research by Papamakarios et al. (Papamakarios et al., 2021). An NF model is a neural network, $g=g_{k}\circ...\circ g_{1}:\mathbb{R}_{z}^{n}\to\mathbb{R}_{\theta}^{n}$, with each layer $g_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, called a flow, that is invertible and differentiable. We let $\theta_{NF}$ denote the parameters of $g$ to be learned, and suppress this notation for simplicity unless needed. Notably, $g$ (and by symmetry, $g^{-1}$) preserves probability densities: if $p_{z}$ is a probability density, then $\theta\mapsto p_{z}(g^{-1}(\theta))|\det J_{g^{-1}}(\theta)|$ is also, which is proven by the change of variables induced by $\theta=g(z)$.
Let $p^{*}_{\theta}$ be an intractable target density over $\mathbb{R}^{n}$ (meaning $p^{*}_{\theta}$ can be computed up to a constant factor, but we cannot sample from $p^{*}_{\theta}$), in our case $p_{\theta}^{*}=p_{ExpM}$. Let $p_{z}$ be a tractable density over $\mathbb{R}^{n}$, esp., $p_{z}$ is Gaussian. Then $g$ can be trained to transform $p_{z}$ into $p_{\theta}$ so that $p_{\theta}\thickapprox p^{*}_{\theta}$ as follows: iteratively sample a batch $z_{i}\sim p_{z},i=1,...,N$ and minimize the empirical KL divergence of the base density $p_{z}$ with the pull-back of the target density, namely $p^{*}_{\theta}(g(z))|\det J_{g}(z)|$. This is “Reverse KL” (RKL) loss, which we approximate with a monte carlo estimate,
$KL[$ $\ p_{z}(z)\ \|\ p^{*}_{\theta}(g(z))|\det J_{g}(z)|\ ]\approxeq$
$\frac{1}{N}\sum_{i=1}^{N}\log\frac{p_{z}(z_{i})}{p^{*}_{\theta}(g(z_{i}))|\det J_{g}(z_{i})|}\ .$
In practice our loss function neglects terms that do not include parameters of $g$ as they do not affect training; in particular, $p^{*}_{\theta}=\exp(f(\theta))/Z$, and we only include the log of the numerator, $\log[p^{*}_{\theta}Z]=\log[\exp(f)]=f$, thereby avoiding computation of the normalization constant $Z$:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\log[p^{*}_{\theta}(g(z_{i}))Z]+\log|\det J_{g}(z_{i})|$ (1)
Once trained, $\theta=g(z)$ with $z\sim p_{z}$ gives approximate samples from previously intractable $p_{\theta}$!
In the ExpM setting $\log(p^{*}_{\theta}(\theta_{i})Z=u(\theta_{i},X,Y)\varepsilon/(2s)$, so performant computation of the utility function is required. Generally, NFs are designed to make computing the log determinant term computationally feasible (e.g., using flows $g_{i}$ so that $J_{g}$ is triangular or that satisfy Theorem B.2) while balancing the expressiveness. In this work, we leverage the Planar Flows of Rezende & Mohamed (Rezende & Mohamed, 2015) and their more expressive generalization, Sylvester Flows (Van Den Berg et al., 2018).
Algorithm 1 Differentially Private ML with Exp+NF

### v4

# Are Normalizing Flows the Key to Unlocking the Exponential Mechanism?

@@
Abstract.
The Exponential Mechanism (ExpM), designed for private optimization, has been historically sidelined from use on continuous sample spaces, as it requires sampling from a generally intractable density, and, to a lesser extent, bounding the sensitivity of the objective function. Any differential privacy (DP) mechanism can be instantiated as ExpM, and ExpM poses an elegant solution for private machine learning (ML) that bypasses inherent inefficiencies of DPSGD. This paper seeks to operationalize ExpM for private optimization and ML by using an auxiliary Normalizing Flow (NF)—an expressive deep network for density learning—to approximately sample from ExpM density. The new method, ExpM+NF, provides an alternative to SGD methods for model training. We prove a sensitivity bound for the $\ell^{2}$ loss function under mild hypotheses permitting ExpM use with any sampling method. To test model training feasibility, we present results on MIMIC-III health data comparing (non-private) SGD, DPSGD, and ExpM+NF training methods’ accuracy and training time. We find that a model sampled from ExpM+NF is nearly as accurate as non-private SGD, more accurate than DPSGD, and ExpM+NF trains faster than Opacus’ state-of-the-art professional DPSGD implementation. Unable to provide a privacy proof for the NF approximation, we present empirical results to investigate privacy including the LiRA membership inference attack of Carlini et al. and the recent privacy auditing lower bound method of Steinke et al. Our findings suggest ExpM+NF provides more privacy than non-private SGD, but not as much as DPSGD, although many attacks are impotent against any model. Ancillary benefits of this work include pushing the state of the art of privacy and accuracy on MIMIC-III healthcare data, exhibiting the use of ExpM+NF for Bayesian inference, showing the limitations of empirical privacy auditing in practice, and providing several privacy theorems applicable to distribution learning. Future directions for proving privacy of ExpM+NF and generally revisiting ExpM with modern sampling techniques are given.
differential privacy, machine learning, density learning, exponential mechanism, privacy auditing, membership inference attack, electronic health records, DPSGD
Vandy Tombs, Robert Bridges co-first authors with equal contributions. Both are corresponding authors.
Special thanks to Cory Watson for making the big machines just work, to Adam Spannaus and Vishakh Gopu for their wisdom with PyTorch, and to Ravi Tandon for insightful conversations.
This manuscript has been co-authored by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the US DOE. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
@@ Abstract.
Figure 1. Figure depicts a to-be-publicized model trained on private data. New technique, ExpM+NF, trains an auxiliary Normalizing Flow (NF) model to approximately sample from the otherwise intractable Exponential Mechanism (ExpM) density to produce near-optimal model parameters $\theta$.
1. Introduction
@@
Winning the 2017 Gödel prize, Differential Privacy (DP) has become an accepted standard for privacy–preserving analytics including machine learning (ML) applications (xiong2014survey, ; ponomareva2023how, ; dajani2017modernization, ; appleDP, ; erlingsson2014rappor, ; ding2017collecting, ). In practice, differentially private ML is achieved almost exclusively by Differentially Private Stochastic Gradient Descent (DPSGD) (abadi2016deep, ; ponomareva2023how, ), and DPSGD is considered the research state of the art as well. To achieve a privacy guarantee, DPSGD clips and adds Gaussian noise to each gradient during SGD training, and upon the conclusion of training, the many privacy guarantees (one per step) are collected into a single, overarching guarantee. Aiding DPSGD’s widespread adoption is 15 years of work on composition theorems (dwork2007sequential, ; dwork2010advanced, ; kairouz2015composition, ; dong2021gaussian, ) which include methods that tighten the privacy bounds for mechanisms that utilize subsampling (e.g., ML training with batches) (abadi2016deep, ; mironov2017renyi, ), and clever techniques that compose privacy curves using the privacy random variable (gopi2021numerical, ).
@@
Yet, in the conventional scenario of training ML on a single private dataset, DPSGD is wasteful by design—privacy is expended so that all gradients may be publicized, while all that is required is the trained model. This results in cumulative privacy degradation with each training step, of which thousands are usually needed. Less problematic, but still not ideal, is that DPSGD can only provide approximate DP, i.e., $\delta>0$, due to relying on composition theorems and on the Gaussian Mechanism, although it does provide “graceful” (continuous) $\varepsilon-\delta$ tradeoffs (see Definition 2.1). Despite 15 years of fruitful research advancing the composition theorems, sub-sampling methods, and implementation techniques, adequate accuracy and privacy have often proved unattainable in practice (googleblog2022, ; suriyakumar2021chasing, ).
Conversely, the Exponential Mechanism (ExpM), developed in 2007 for private optimization (mcsherry2007mechanism, ), has been historically sidelined from ML training by two obstructions. Primarily, sampling from the required distribution is computationally infeasible except for finite or simple domains, reducing ExpM to multinomial sampling (friedman2010data, ; kapralov2013differentially, ; vinterbo2012differentially, ). Notwithstanding these examples, ExpM is historically intractable for most ML training and more generally any application involving a high–dimensional continuous sample space (kasiviswanathan2011what, ; ji2014differential, ; minami2016differential, ; near2022programming, ). Secondly, a sensitivity bound must be proven for the given utility (loss) function, which is difficult in general, e.g., (minami2016differential, ). Notably, ExpM furnishes strong DP ($\delta=0$) and has a proven guarantee on its utility (accuracy) (mcsherry2007mechanism, ). Further, sampling a model from ExpM bypasses the incremental decline in privacy suffered by DPSGD; hence, ExpM’s application to differentially private ML is potentially transformational, not just in method but in the privacy-accuracy tradeoff afforded. Additoinally, it is known that any DP mechanism can be instantiated as ExpM (mcsherry2007mechanism, ). In short, work to unlock ExpM promises large impact for differentially privacy, especially in ML application, and hence, research to discover how to operationalize ExpM is due.
Meanwhile, sampling methods have progressed independently. In particular, Normalizing Flows (NFs), expressive deep networks for approximating intractable distributions, have emerged as alternatives to Markov Chain Monte Carlo (MCMC) techniques that provide rapid sampling once trained, bypass inefficiencies of MCMC, and are generally effective on high–dimensional settings (gabrie2021efficient, ; papamakarios2021normalizing, ). Hence the pairing of ExpM with an NF as a sampling methodology is a now-available solution that can potentially unlock the ExpM!
In this paper, we introduce ExpM+NF, in particular, as a general method of model training, depicted in Figure 1. Instead of an SGD scheme, we leverage ExpM to optimize the loss function, and employ an auxiliary NF model for sampling (ExpM+NF). As our results show, it is an effective method providing very accurate results, but we are unable to provide a proof of privacy for the NF approximation of ExpM’s density. By design ExpM+NF would eliminate the iterative privacy loss observed in DPSGD and offers potential for a strong, user-specified privacy guarantee ($\varepsilon>0,\delta=0$) with near optimal parameters.
Below, we document our research to explore if and exactly how one may train an ML model via ExpM+NF; we examine its computational expense, accuracy, and privacy. We include positive and negative results, as well as identified limitations and potentially fruitful future directions. Overall, we believe the idea of using modern sampling methods with ExpM has the potential to transform differentially private ML. Our hope is that the directions put forth in this document are fodder for this research community to identify a successful path in this regard.
1.1. Contributions and Findings
Contribution 1: We prove a sensitivity bound for $\ell^{2}$ (equiv. MSE) loss for supervised learning and regression tasks with bounded targets (Theorem 3.1). This accommodates classification tasks via binary/one–hot encoding, permitting use of ExpM with any sampling technique.
@@ 1.1. Contributions and Findings
Contribution 2.b: An ancillary benefit of the previous experiment is that our DPSGD implementation is the state of the art (in privacy and accuracy) on the MIMIC-III ICU binary prediction benchmarks. See Appendix E.2.
Contribution 2.c: Using ExpM+NF is essentially training the NF to approximate a posterior distribution. In the non-private scenario, ExpM+NF facilitates Bayesian inference and uncertainty quantification (UQ) as it allows rapid sampling of parameters from the posterior. A simple example is exhibited, see Appendix F.
Contribution 3: Just as accuracy on a held-out test set is used to gauge accuracy, we implement membership inference attacks (MIA) against (non-private) SGD, DPSGD, and ExpM+NF to gain empirical results on the methods’ privacy. Using the (state of the art MIA) Likelihood Ratio Attack (LiRA) (carlini2022membership, ) our conclusions vary per model. When using the LR model, we find that DPSGD is essentially unattackable, while ExpM+NF is less resilient than DPSGD but more so than non-private SGD. Our LiRA results for the GRU-D model are inconclusive as all models are essentially unattackable. In short, the evidence is that ExpM+NF provides some privacy, but not as much as DPSGD for the same $\tilde{\varepsilon}$ parameter, yet it provides better accuracy for the same $\tilde{\varepsilon}$.
Contribution 4: Recent exciting work by Steinke et al. (steinke2024privacy, ) proves a lower bound on the privacy guarantee $\varepsilon$, that holds with known likelihood, and is computable from the accuracy of a black-box attack seeking to identify known “canaries” in the training set. We apply this privacy auditing methodology, but find that in all cases the lower bound is 0. Our results provide a glimpse at the difficulties with privacy auditing in practice—if a model is resistant to the attack used by Steinke et al., then the auditing methodology is of little help.
Contribution 5: Our final contributions are mathematical advancements towards a formal differential privacy proof for ExpM+NF. While privacy of ExpM+NF is not proven, we present more widely applicable math results, and future methods of proof as well as obstructions identified. Specific advancements include the following: We define a metric $d_{DP}$ on probability densities so that convergence in this metric preserves $\varepsilon$-DP, and show it is strictly stronger than uniform convergence, which can provide ($\varepsilon,\delta>0$)-DP under appropriate hypothesis. (This shows the need for the new metric). We provide a Privacy Squeeze Theorem for proving privacy of a mechanism whose density is sandwiched between two densities of known privacy and similarly a privacy Projection Theorem.
1.2. Limitations
There are three primary limitations with ExpM+NF, currently. First and foremost is the lack of privacy proof for ExpM+NF. (Interestingly, works on MCMC sampling methods with ExpM (lin2024tractable, ; gopi2022private, ) provide privacy proofs with no implementation or testing, so feasibility of the methods are unknown). The primary obstruction and need is a proof that for any training dataset $X\in D$ (in our universe of datasets), the trained NF must lie either close to the target density or in some relationship to densities that have known privacy bounds. Theorems and formal definitions to make these notions rigorous are the topic of Section 7, along with future directions for research. Second, ExpM+NF requires a sensitivity bound for each loss function, and thus far we have only provided such a proof for $\ell^{2}$ (equiv. mean squared error) loss; while this does not prove limiting in our experiments, future research is required for application to other loss functions. By comparison, DPSGD is readily applicable to theoretically any loss function. Third, ExpM+NF requires a completely different training regime (training an auxiliary NF model whose loss function depends the desired model and its loss function); hence, custom software engineering is currently required for use on each model architecture (e.g., on LR, GRU-D, …).
We also identify experimental limitations. Hyperparameter searching was required for all methods and we ignored privacy of hyperparameter searching which is not reealistic. This choice was intentional—we aim to compare the very best DPSGD and ExpM+NF can perform. The limited abilities of both the empirical privacy auditing experiments (see Section 6.1 containing results of the LiRA membership inference attack of Carlini et al. and black–box attack of Steinke et al.) narrow the aperture for empirically investigating ExpM+NF’s privacy. Aside from impotence of these attacks, a degredation of defenses as the $\varepsilon$ increased is expected for both ExpM+NF and DPSGD and is not observed. This may be caused in part by the size and any obfuscation (purposefully before release by the curators, or inadvertently in the benchmark preparation pipeline, e.g., imputation of missing results) performed on the MIMIC-III data. Optimistically, the poor results of these attacks are good news for the health data community!
2. Background
[…]
2.1. Differential Privacy Background
@@
A mechanism $f:\Theta\times D\to\mathbb{R}$ satisfies $(\varepsilon,\delta)$-Differential Privacy or is $(\varepsilon,\delta)$-DP if and only if $P(f(\theta,X)\in A)\leq P(f(\theta,X^{\prime})\in A)e^{\varepsilon}+\delta$ for all neighboring $X,X^{\prime}\in D,\theta\in\Theta,\varepsilon\geq 0$ and for all measurable sets $A$. When $\delta=0$, we say $f$ is $\varepsilon$-DP.
The $\varepsilon$ term is an upper bound on the privacy loss. Approximate DP, or $(\varepsilon,\delta)$-DP, provides a relaxed definition—$\delta$ is the probability that the mechanism $f$ is not guaranteed to be $\varepsilon$-DP. This “probability of failure” interpretation has led many to suggest that $\delta$ should be no larger than $1/|X|$ (dwork2014algorithmic, ).
To achieve $(\varepsilon,\delta)$-DP in practice, one calibrates the randomness (mechanism’s variance) by how sensitive the output of the mechanism is across all neighboring datasets.
@@
The sensitivity of a mechanism $f:\Theta\times D\to\mathbb{R}$ is $s=\sup|f(\theta,X)-f(\theta,X^{\prime})|$ where the supremum is over neighboring $X,X^{\prime}\in D$, and all $\theta\in\Theta$. (If multiple mechanisms are present, say $f_{1},f_{2}$ we will use $\Delta(f_{1}),\Delta(f_{2})$ for their respective sensitivities rather than $s$.
Definition 2.3 (ExpM).
For utility function $u:\Theta\times D\to\mathbb{R}$ with sensitivity $s$, the Exponential Mechanism (ExpM) returns $\theta\in\Theta$ with likelihood $p_{ExpM}(\theta;X)\propto\exp(\varepsilon u(\theta,X)/(2s))$.
Theorem 2.4 (ExpM satisfies $\varepsilon$-DP (mcsherry2007mechanism, ; dwork2014algorithmic, )).
@@
The original treatment of ExpM (mcsherry2007mechanism, ) furnishes utility theorems—rigorous guarantees that bound how far ExpM samples can be from the optima.
2.2. Normalizing Flows (NFs)
We refer readers to the survey of Papamakarios et al. (papamakarios2021normalizing, ). An NF model is a neural network, $g=g_{k}\circ...\circ g_{1}:\mathbb{R}_{z}^{n}\to\mathbb{R}_{\theta}^{n}$, with each layer $g_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$, called a flow, that is invertible and differentiable. We let $\beta_{\text{NF}}$ denote the parameters of $g$ to be learned, and suppress this notation for simplicity unless needed. Notably, $g$ (and by symmetry, $g^{-1}$) preserves probability densities: if $p_{z}$ is a probability density, then $\theta\mapsto p_{z}(g^{-1}(\theta))|\det J_{g^{-1}}(\theta)|$ is also.
Suppose the target density, $p^{*}_{\theta}$ can be computed up to a constant factor, but we cannot sample from $p^{*}_{\theta}$. Let the base density $p_{z}$ be a tractable density over $\mathbb{R}^{n}$. Then $g$ can be trained to transform $p_{z}$ into $p_{\theta}$ so that $p_{\theta}\thickapprox p^{*}_{\theta}$ as follows: iteratively sample a batch $z_{i}\sim p_{z},i=1,...,N$ and minimize the empirical KL divergence of the base density $p_{z}$ with the pull-back of the target density, namely $p^{*}_{\theta}(g(z))|\det J_{g}(z)|$. This is the “Reverse KL” (RKL) loss, which we approximate with a monte carlo estimate,
$KL[$ $\ p_{z}(z)\ \|\ p^{*}_{\theta}(g(z))|\det J_{g}(z)|\ ]\approxeq\frac{1}{N}\sum_{i=1}^{N}\log\frac{p_{z}(z_{i})}{p^{*}_{\theta}(g(z_{i}))|\det J_{g}(z_{i})|}\ .$
In practice our loss function neglects $\log{p_{z}(z_{i})}$ since it does not include parameters of $g$. Further, $\log{p^{*}_{\theta}(g(z_{i}))|\det J_{g}(z_{i})|}$ is only required up to a constant allowing avoidance of intractable normalization terms, e.g., when $p^{*}(\theta)=e^{u(\theta)}/Z$. Once trained, $\theta=g(z)$ with $z\sim p_{z}$ gives approximate samples from previously intractable $p_{\theta}$!
In this work, we leverage the Planar Flows of Rezende & Mohamed (rezende2015variational, ) and their more expressive generalization, Sylvester Flows (berg2018sylvester, ); the target is $p_{\theta}^{*}=p_{ExpM}$; and base, $p_{z}$, is Gaussian.